OpenStudy (anonymous):

Evaluate the integral by making the given substitution. integral of dx/((5x+14)^2)

OpenStudy (anonymous):

u=5x+14, du=5dx, du/5=dx so substitute this into dx/((5x+14)^2) to get this into u so integral of dx/((5x+14)^2) = (1/5) integral of dx/(u^2) = (1/5) (-1/u) = (-1/5u) now substitute u=5x+14 into this and you get y=-1/(25x+60) as the final answer

OpenStudy (anonymous):

haha it is 70...not 60

OpenStudy (anonymous):

so its my mistake

OpenStudy (anonymous):

i mean 25x+70 is right but i forgot the negative sign Thanks i had a mistake too

OpenStudy (anonymous):

no worries...